# Provably Efficient RLHF Pipeline: A Unified View from Contextual Bandits

### v1

@@ Abstract
Reinforcement Learning from Human Feedback (RLHF) is a widely used approach for aligning Large Language Models (LLMs) with human preferences. While recent advancements have provided valuable insights into various stages and settings of RLHF, a comprehensive theoretical understanding of the entire RLHF pipeline remains lacking. Towards this end, we propose a unified framework for the RLHF pipeline from the view of contextual bandits and provide provable efficiency guarantees. In particular, we decompose the RLHF process into two distinct stages: (post-)training and deployment, exploring both passive and active data collection strategies during the training phase. By employing the Bradley-Terry preference model with a linearly parameterized reward function, we reformulate RLHF as a contextual preference bandit problem. We then develop novel algorithms for each stage, demonstrating significant improvements over existing approaches in both statistical and computational efficiency. Finally, we apply our method to train and deploy Llama-3-8B-Instruct on the Ultrafeedback-binarized dataset, and empirical results confirm the effectiveness of our approach.
@@ 1 Introduction
Table 1: Comparison between previous works and approach in terms of the statistical and computational efficiency across different stages of the RLHF pipeline. Our pipeline consists of two stages: training and deployment, where we explore both passive and active data collection strategies in the training stage. The statistical efficiency is measured using the sub-optimality gap (for the training stage) and regret (for the deployment stage), while the computational efficiency is evaluated based on time and storage complexity. Here, $d$ is the dimension of the feature space, $T$ is the number of samples, $\kappa$ is the non-linearity coefficient which can be exponentially large, $\mathbb{E}_{x\sim\rho}\left(\phi(x,\pi^{*}(x))-v\right)$ is the concentratability vector, $V_{T}$ and $\mathcal{H}_{T}$ are two kinds of local norm that satisfies $\mathcal{H}_{T}\succeq(1/\kappa)V_{T}$.
Stage Data Reference Sub-optimality Gap/Regret Time Storage
Training Passive Zhu et al. (2023) $\widetilde{\mathcal{O}}\Big{(}\kappa\sqrt{d}\cdot\big{\lVert}\mathbb{E}_{x\sim\rho}\left(\phi(x,\pi^{*}(x))-v\right)\big{\rVert}_{V_{T}^{-1}}\Big{)}$ $\mathcal{O}(T\log T)$ $\mathcal{O}(T)$
Ours (Theorem 1) $\widetilde{\mathcal{O}}\Big{(}\sqrt{d}\cdot\big{\lVert}\mathbb{E}_{x\sim\rho}\left(\phi(x,\pi^{*}(x))-v\right)\big{\rVert}_{\mathcal{H}_{T}^{-1}}\Big{)}$ $\mathcal{O}(T)$ $\mathcal{O}(T)$
Active Das et al. (2024) $\widetilde{\mathcal{O}}\big{(}d\sqrt{\kappa/T}\big{)}$ $\mathcal{O}(T^{2}\log T)$ $\mathcal{O}(T)$
Ours (Theorem 2) $\widetilde{\mathcal{O}}\big{(}d\sqrt{\kappa/T}\big{)}$ $\mathcal{O}(T)$ $\mathcal{O}(1)$
Deployment Online Saha (2021) $\widetilde{\mathcal{O}}\big{(}d\kappa\sqrt{T}\big{)}$ $\mathcal{O}(T^{2}\log T)$ $\mathcal{O}(T)$
Ours (Theorem 3) $\widetilde{\mathcal{O}}\big{(}d\sqrt{\kappa T}\big{)}$ $\mathcal{O}(T)$ $\mathcal{O}(1)$
@@
Reinforcement Learning from Human Feedback is a key technique for training large language models using human feedback (Ouyang et al., 2022; Bai et al., 2022). The RLHF process involves collecting data, each consisting of a prompt, a pair of responses, and a human preference label indicating the preferred response. Then, a reward model is trained to predict human preferences, and the LLM is fine-tuned based on the reward model by RL algorithms, such as PPO (Schulman et al., 2017).
Given the notable success of RLHF, recent efforts have been devoted to developing a deeper theoretical understanding of this approach. Zhu et al. (2023) investigated the standard offline setting, where the learner is provided with a fixed dataset and aims to learn a policy that maximizes the expected reward. In this setting, since the learner has no control over the data collection process, the quality of the dataset becomes crucial to the performance of the learned policy and the resulting policy often performs poorly when faced with out-of-distribution data (Burns et al., 2024). In practice, the Claude (Bai et al., 2022) and LLaMA-2 (Touvron et al., 2023) projects have demonstrated that iterative RLHF can significantly enhance model performance. Motivated by these practical advancements, Xiong et al. (2024) investigated the iterative setting, where the learner randomly selects prompts from the candidate set, generates response pairs by sampling from the current policy, and then queries for feedback to update the policy accordingly. They showed that RLHF benefits from this iterative data collection process. While uniform prompt sampling as a simple approach has been proven effective, Das et al. (2024) demonstrated that it can lead to inefficient sampling and the sub-optimality gap may fail to decrease as the number of samples increases. To this end, they studied the active setting where the learner is allowed to select both prompts and responses, query for human feedback, and update the policy accordingly.
Though these works have provided valuable insights into various stages and settings of RLHF, a comprehensive theoretical understanding of the entire RLHF pipeline remains lacking. Towards this end, in this work, we propose a unified framework for the RLHF pipeline from the view of contextual bandits. Specifically, building on recent theoretical advancements (Zhu et al., 2023; Xiong et al., 2024), we leverage the Bradley-Terry preference model with a linearly parameterized reward function to reformulate RLHF as a contextual bandit problem (Abe and Long, 1999). We then decompose the RLHF process into two distinct stages: (post-)training and deployment. In the training stage, the learner’s goal is to develop a high-performing model with minimal sample queries and computational resources, where we explore both passive and active data collection strategies. In the deployment stage, the trained model is used to select good responses to user inputs while concurrently gathering informative feedback from the user for further refinement.
Building on this unified view, we revisit the existing theoretical advancements in RLHF and observe that they primarily focus on specific stages and settings of the overall pipeline. Consequently, an understanding of the entire pipeline remains lacking. Furthermore, while they have proposed various algorithms with theoretical guarantees for different stages of RLHF, the high statistical and computational complexity of their methods presents a significant challenge for practical applications. Thus, a natural question arises:
Can we develop provably efficient algorithms with improved statistical and computational efficiency for the RLHF pipeline?
In this work, we answer this question affirmatively by proposing novel algorithms for each stage of the RLHF pipeline, improving both statistical and computational efficiency over existing methods. Specifically, our contributions are summarized as follows:
  • We propose a unified framework for the RLHF pipeline from the perspective of contextual bandits. We decompose the RLHF process into two stages: (post-)training and deployment, and explore both passive and active data collection strategies during training. This perspective allows us to view several prior works as focusing on specific components of the pipeline.
  • We propose algorithms that are computationally and statistically more efficient than existing methods for each stage of the RLHF pipeline. Specifically, we employ online mirror descent with a tailored local norm to replace the standard maximum likelihood estimation. Additionally, we design distinct query criteria for different stages. Table 1 summarizes the comparison between previous works and ours about statistical and computational efficiency.
  • We train and deploy Llama-3-8B-Instruct on the Ultrafeedback-binarized dataset using our method, testing both passive and active data collection strategies in the training stage. The empirical results validate our theoretical findings and confirm the effectiveness of our method.
Organization.
The remainder of this paper is organized as follows. We review related work in Section 2 and introduce the problem setup in Section 3. Our pipeline is presented in Sections 4 and 5, focusing on the training and deployment stages, respectively. We present experiments in Section 6 and conclude in Section 7. All proofs and experiment details are deferred to the appendix.
Notations.
For a vector $x\in\mathbb{R}^{d}$ and a positive semi-definite matrix $A\in\mathbb{R}^{d\times d}$, we define $\lVert x\rVert_{A}=\sqrt{x^{\top}Ax}$. We use the notation $\mathcal{O}(\cdot)$ to highlight various dependencies depending on the context. For the statistical efficiency, $\mathcal{O}(\cdot)$ omits only constant factors. For computational efficiency, $\mathcal{O}(\cdot)$ solely emphasizes the dependence on the number of samples as this is the primary factor influencing the complexity. Additionally, we employ $\widetilde{\mathcal{O}}(\cdot)$ to hide polylogarithmic factors.
@@ 2 Related Work
In this section, we review the works most closely related to ours.
Reinforcement Learning from Human Feedback.
Learning from human preferences with deep learning models dates back to Christiano et al. (2017) and has been recently popularized by the success of large language models (Bai et al., 2022; Touvron et al., 2023; Achiam et al., 2023). Existing results generally fall into two categories: reward-based and reward-free. Reward-based methods typically consist of two steps: reward modeling and reward maximization (Ouyang et al., 2022). Given preference data, the reward model is trained to predict preference labels, and the LLM is fine-tuned based on the reward model using RL algorithms such as PPO (Schulman et al., 2017). In contrast, reward-free methods directly optimize the LLM using human preference feedback, without relying on explicit reward modeling (Rafailov et al., 2023; Azar et al., 2024). Among these works, Direct Preference Optimization (DPO) (Rafailov et al., 2023) is one of the most popular reward-free methods, which treats generative models directly as reward models and optimizes them using human preference feedback. In this work, we focus on the reward-based RLHF methods.
Given the empirical success of RLHF, recent efforts have been devoted to developing a deeper theoretical understanding of this approach. Zhu et al. (2023) investigated the standard offline setting, where the learner is provided with a fixed dataset and aims to learn a policy that maximizes the expected reward. Xiong et al. (2024) investigated the iterative setting, where the learner randomly selects prompts from the candidate set, generates response pairs by sampling from the current policy, and then queries for feedback to update the policy accordingly. Das et al. (2024) studied the active setting where the learner is allowed to select both prompts and responses, query for feedback, and update the policy accordingly. Liu et al. (2024) focused on the dual active setting where the learner is allowed to select conversations and teachers simultaneously for feedback. Essentially, these works focus on specific stages and settings within the RLHF pipeline, while we offer a unified view on the entire process. Furthermore, we propose algorithms that are both computationally and statistically more efficient than existing methods for each stage of the RLHF pipeline.
@@
Dueling bandits is a variant of the multi-armed bandit problem that incorporates pairwise comparison feedback (Yue et al., 2012). Contextual dueling bandits further extend this framework by integrating context information (Dudík et al., 2015). In this setting, Saha (2021) studied the $K$-arm contextual dueling bandit problem and proposed an algorithm with a near-optimal regret guarantee. Saha et al. (2023) extended the contextual dueling bandit problem to the reinforcement learning setting, presenting an algorithm with a sublinear regret bound. Sekhari et al. (2024) further studied the contextual dueling bandit problem with active learning, where the learner is allowed to query for feedback adaptively and the goal is to minimize the regret while minimizing the number of queries simultaneously. However, these works primarily focus on the online setting, where the context is provided in an arbitrary and online manner and goal is to minimize cumulative regret. In this work, we consider the entire RLHF pipeline, including both training and deployment stages, and explore both passive and active data collection strategies in the training stage.
@@
Active learning is a well-established field that aims to reduce the labeling cost by selecting the most informative samples for annotation (Settles, 2009). The first line of work focuses on the pool-based setting (Seung et al., 1992; Freund et al., 1997; Huang et al., 2010). In this setting, the learner strategically selects a batch of the most informative data at each step and queries labels only for this selected batch. The learner then uses this labeled batch to update the model. Once the model is enhanced, the learner selects another batch of data and queries labels for it, continuing the iterative training process. The other line of work focuses on the online active setting (Cesa-Bianchi et al., 2004, 2006). In this setting, the learner sequentially observes data points and decides whether to query labels for them. Ji et al. (2024) explored the online active RLHF setting where prompts are provided in an arbitrary and online manner, with the goal of minimizing cumulative regret while minimizing the number of queries simultaneously. In this work, we leverage the pool-based active learning strategy in the training stage of the RLHF pipeline, where the learner is allowed to select the dataset actively.
@@ 3 Problem Setup
In this section, we introduce the problem setup for RLHF.
Following recent advancements in RLHF (Zhu et al., 2023; Xiong et al., 2024), we formulate RLHF as a contextual bandit problem. Specifically, we have a set of contexts $\mathcal{X}$ and a set of possible actions $\mathcal{A}$ per context. To learn with human preference feedback, the learner selects a tuple $(x,a,a^{\prime})$ to present to the human, where $x\in\mathcal{X}$ is the context, $a,a^{\prime}\in\mathcal{A}$ are the actions. The human then provides a binary preference feedback $y\in\{0,1\}$, where $y=1$ indicates the human prefers action $a$ over action $a^{\prime}$, and $y=0$ otherwise. We study the commonly used Bradley-Terry (BT) model in preference learning (Bradley and Terry, 1952), which assumes that the human’s preference is generated by a logistic function of the difference in the rewards of the two actions.
@@
$\mathbb{P}\left[y=1\mid x,a,a^{\prime}\right]=\frac{\exp\left(r(x,a)\right)}{\exp\left(r(x,a)\right)+\exp\left(r\left(x,a^{\prime}\right)\right)},$
where $r$ is a latent reward function.
We consider the linear reward model which is widely used in the RLHF literature (Zhu et al., 2023; Ji et al., 2024; Das et al., 2024).
@@
It holds that $r(x,a)=\phi(x,a)^{\top}\theta^{*}$ where $\phi(x,a):\mathcal{X}\times\mathcal{A}\to\mathbb{R}^{d}$ is the known and fixed feature map, and $\theta^{*}\in\mathbb{R}^{d}$ is the unknown parameter vector. Furthermore, we assume $\lVert\phi(x,a)\rVert_{2}\leq L$ for all $x\in\mathcal{X}$ and $a\in\mathcal{A}$ and $\theta^{*}\in\Theta$ where $\Theta=\{\theta\in\mathbb{R}^{d}\mid\langle\bm{1},\theta\rangle=0,\|\theta\|_{2}\leq B\}$.
@@
Based on the linear model, we can rewrite the probability as
$\mathbb{P}\left[y=1\mid x,a,a^{\prime}\right]=\sigma\left(\phi(x,a)^{\top}\theta^{*}-\phi(x,a^{\prime})^{\top}\theta^{*}\right),$ (1)
where $\sigma(w)=\frac{1}{1+\exp(-w)}$ is the sigmoid function. Next, we introduce a key quantity that captures learning complexity under the BT model.

### v2

# Provably Efficient Online RLHF with One-Pass Reward Modeling

@@ Abstract
Reinforcement Learning from Human Feedback (RLHF) has shown remarkable success in aligning Large Language Models (LLMs) with human preferences. Traditional RLHF approaches rely on a fixed dataset, which often suffers from limited coverage. To this end, online RLHF has emerged as a promising direction, enabling iterative data collection and model improvement. Despite its potential, this paradigm faces a key bottleneck: the requirement to continuously integrate new data into the historical dataset and re-optimize the model from scratch at each iteration, resulting in computational and storage costs that grow linearly with the number of iterations. In this work, we address this challenge by proposing a one-pass reward modeling method that does not require storing the historical data and can be computed in constant time. Specifically, we first formalize RLHF as a contextual preference bandit problem and design an online mirror descent algorithm with a tailored local norm to replace the standard maximum likelihood estimation for reward modeling. We then apply our method to various online RLHF settings, including passive data collection, active data collection, and deployment-time adaptation. We provide theoretical guarantees showing that our method improves both statistical and computational efficiency. Finally, we provide practical algorithms and conduct experiments using Llama-3-8B-Instruct and Qwen2.5-7B-Instruct models on the Ultrafeedback-binarized and Mixture2 datasets, validating the effectiveness of our proposed method.
@@ 1 Introduction
Reinforcement Learning from Human Feedback is a critical technique for training large language models using human preference feedback (Ouyang et al., 2022; Bai et al., 2022). Typical RLHF methods involve collecting extensive data, each consisting of a prompt, a pair of responses, and a preference label indicating which response is preferred. Then, a reward model is trained to predict the human preference, and the LLM is fine-tuned based on the reward model by the RL algorithms.
Traditional RLHF methods primarily rely on fixed preference datasets, which typically suffer from limited coverage. As a result, the learned reward models struggle to generalize to out-of-distribution (OOD) samples, constraining the effectiveness of the aligned models. To address this, online RLHF has emerged as a promising paradigm, enabling iterative data collection and model improvement. The general process can be described as (i) collect the preference data; (ii) update the model using the collected preference data. The above two steps are repeated for several iterations to boost model performance. In practice, the Claude (Bai et al., 2022) and LLaMA-2 (Touvron et al., 2023) projects have demonstrated that online RLHF can significantly enhance model performance.
Despite its empirical success, online RLHF introduces significant computational challenges. Specifically, the typical process of online RLHF involves continuously integrating newly collected data into the dataset and re-optimizing the model from scratch over the expanded dataset. While this strategy is statistically efficient, its computational and storage costs scale linearly with the number of iterations, which becomes prohibitive in long-term iterations, especially on edge devices where computation and memory resources are inherently limited. This raises a pressing question:
Can we design online RLHF algorithms that are both statistically and computationally efficient?
[…]
Setting Context Action Gap/Regret Time Storage Reference
[…]
$\widetilde{\mathcal{O}}\big{(}\sqrt{d}\cdot\lVert\Phi\rVert_{\mathcal{H}_{T}^{-1}}\big{)}$ $\mathcal{O}(1)$ $\mathcal{O}(1)$ Ours (Theorem 1)
[…]
$\widetilde{\mathcal{O}}\big{(}d\sqrt{\kappa/T}\big{)}$ $\mathcal{O}(1)$ $\mathcal{O}(1)$ Ours (Theorem 2)
[…]
$\widetilde{\mathcal{O}}\big{(}d\sqrt{\kappa T}\big{)}$ $\mathcal{O}(1)$ $\mathcal{O}(1)$ Ours (Theorem 3)
In this work, we provide an affirmative answer to this question in the setting of contextual preference bandits with linearly parameterized reward functions. Specifically, building on recent theoretical advancements in RLHF (Zhu et al., 2023; Das et al., 2024; Ji et al., 2024), we formulate the RLHF problem as a contextual dueling bandit problem (Yue et al., 2012; Saha, 2021). While prior work has explored this formulation, most existing methods focus on statistical efficiency and overlook the growing computational burden. To bridge this gap, we introduce a novel one-pass reward modeling algorithm which does not require storing the historical data and can be computed in constant time. Our method is based on the online mirror descent framework with a tailored local norm that captures second-order information. We then apply our method to several online RLHF settings, including passive data collection, active data collection, and deployment-time adaptation. We establish theoretical guarantees showing that our method improves both statistical and computational efficiency. Table 1 summarizes the comparison of our method with the existing methods.
To enable usage in LLMs, we develop practical variants of our method. Direct computation and storage of the Hessian matrix is prohibitively expensive; thus, we propose an efficient approximation using Hessian-Vector Products (HVP) combined with conjugate gradient descent, avoiding explicit second-order information and relying only on first-order computation. Additionally, we employ rejection sampling to approximate model uncertainty in a computationally efficient manner. With the above techniques, we conduct experiments using the LLaMA-3-8B-Instruct (Llama Team, 2023) and Qwen2.5-7B-Instruct (Qwen Team, 2024) models on the Ultrafeedback-binarized (Cui et al., 2023) and Mixture2 (Dong et al., 2024) datasets, validating the effectiveness of our method.
To summarize, our contributions are as follows:
  • By formulating the RLHF problem as a contextual dueling bandit, we propose a novel one-pass reward modeling algorithm and establish the corresponding estimation error bound. Our method is built upon the online mirror descent framework and incorporates a carefully designed local norm that captures second-order information for improved learning efficiency.
  • We apply our method to a broad range of online RLHF settings, including passive data collection, active data collection, and deployment-time adaptation. For each setting, we design tailored algorithms and establish corresponding theoretical guarantees, demonstrating that our approach achieves improved statistical and computational efficiency than existing methods.
  • We develop practical algorithms by approximating the update using Hessian-Vector Products combined with conjugate gradient descent, and estimating uncertainty via rejection sampling. Based on the above techniques, we conduct empirical evaluations using LLaMA-3-8B-Instruct and Qwen2.5-7B-Instruct models on Ultrafeedback-binarized and Mixture2 datasets, showing our method achieves better performance in both statistical and computational efficiency.
Organization. The rest of the paper is organized as follows: Section 3 introduces the problem setup. Section 4 presents our proposed one-pass reward modeling method. Section 5 applies our method to different online RLHF settings. Section 6 provides the practical implementations. Section 7 presents experimental results. Section 8 concludes the paper. We defer the proofs, experiment details, and more experimental results to the appendix.
@@ 2 Related Work
In this section, we review the works most closely related to ours, including offline RLHF, online RLHF, contextual dueling bandits, and active learning.
Offline RLHF.
Learning from human preferences with deep learning models dates back to Christiano et al. (2017) and has been recently popularized by the success of large language models (Bai et al., 2022; Touvron et al., 2023; Achiam et al., 2023). Existing results generally fall into two categories: reward-based and reward-free. Reward-based methods typically consist of two steps: reward modeling and reward maximization (Ouyang et al., 2022). Given preference data, the reward model is trained to predict preference labels, and the LLM is fine-tuned based on the reward model using RL algorithms such as PPO (Schulman et al., 2017). In contrast, reward-free methods directly optimize the LLM using human preference feedback, without relying on explicit reward modeling (Rafailov et al., 2023; Azar et al., 2024). Among these works, Direct Preference Optimization (DPO) (Rafailov et al., 2023) is one of the most popular reward-free methods, which treats generative models directly as reward models and optimizes them using human preference feedback. Given the empirical success of RLHF, recent efforts have been devoted to developing a deeper theoretical understanding of this approach. Zhu et al. (2023) proposed to formulate RLHF as a contextual preference bandit problem and proved the maximum likelihood estimator converges under both the Bradley-Terry (BT) model and Plackett-Luce (PL) model with linear reward functions.
Online RLHF.
Offline RLHF methods rely on fixed preference datasets, which often suffer from limited data coverage. Consequently, the resulting reward models struggle to generalize to out-of-distribution samples, thereby limiting the effectiveness of the aligned models. To overcome this limitation,online RLHF has emerged as a promising alternative, enabling iterative data collection and continuous model refinement. The works (Dong et al., 2023; Guo et al., 2024; Yuan et al., 2024; Wu et al., 2025) have demonstrated that online iterative variants of direct preference learning algorithms significantly outperform their offline counterparts. Xiong et al. (2024) identified key challenges in offline RLHF and theoretically demonstrated the potential benefits of online exploration. Recent work has incorporated optimism-driven bonus terms into the DPO loss to encourage exploration in online RLHF (Xie et al., 2025; Cen et al., 2025; Zhang et al., 2025). These studies primarily focus on the sample efficiency, but do not consider the accompanying increase in computational complexity. In this paper, we propose algorithms that achieve both statistical and computational efficiency.
@@
Dueling bandits are a variant of the multi-armed bandit problem in which the learner sequentially selects a pair of arms and receives noisy binary feedback (Yue et al., 2012). The contextual dueling bandit framework extends this setting by incorporating contextual information (Dudík et al., 2015; Saha, 2021; Bengs et al., 2022). Within this framework, Saha (2021) studied the $K$-armed contextual dueling bandit problem, and Saha et al. (2023) further extended it to the reinforcement learning setting. Additionally, Sekhari et al. (2024) investigated the contextual dueling bandit problem under an active learning paradigm, where the learner adaptively queries for feedback, aiming to minimize both regret and the number of queries. To move beyond linear reward functions, Verma et al. (2025) introduced the neural dueling bandit problem, modeling the reward function using neural networks. These prior works commonly rely on maximum likelihood estimation to learn the reward function, leading to computational complexity that grows linearly with the number of iterations. In contrast, we propose algorithms that maintain constant per-iteration computational complexity, while preserving statistical efficiency.
@@
Active learning is a paradigm that aims to reduce the labeling cost by selecting the most informative samples for annotation (Settles, 2009). In general, existing work can be categorized into two settings: pool-based and stream-based. The pool-based setting (Seung et al., 1992; Freund et al., 1997; Huang et al., 2010) involves the learner iteratively selecting a batch of informative samples from a large unlabeled pool, querying their labels, updating the model, and repeating this process. In contrast, the stream-based setting (Cesa-Bianchi et al., 2004, 2006; Cacciarelli and Kulahci, 2024) requires the learner to sequentially observe data points and decide in real time whether to query their labels. Within the context of RLHF, Das et al. (2024) studied pool-based active learning, while Ji et al. (2024) focused on the stream-based setting. In this work, we focus on the pool-based strategy, which can be naturally extended to the stream-based scenario.
@@ 3 Problem Setup
Following recent advancements in RLHF (Zhu et al., 2023; Das et al., 2024; Xiong et al., 2024), we formulate RLHF as a contextual bandit problem. Specifically, we have a set of contexts $\mathcal{X}$ and a set of possible actions $\mathcal{A}$ per context. To learn with human preference feedback, the learner selects a tuple $(x,a,a^{\prime})$ to present to the human, where $x\in\mathcal{X}$ is the context, $a,a^{\prime}\in\mathcal{A}$ are the actions. The human then provides a binary preference feedback $y\in\{0,1\}$, where $y=1$ indicates that the human prefers action $a$ over action $a^{\prime}$, and $y=0$ otherwise. We study the commonly used Bradley-Terry (BT) model in preference learning (Bradley and Terry, 1952), which assumes that the human’s preference is generated by a logistic function of the difference in the rewards of the two actions.
@@
$\mathbb{P}\left[y=1\mid x,a,a^{\prime}\right]=\frac{\exp\left(r(x,a)\right)}{\exp\left(r(x,a)\right)+\exp\left(r\left(x,a^{\prime}\right)\right)},$ (1)
where $r:\mathcal{X}\times\mathcal{A}\to\mathbb{R}$ is a latent reward function.
We consider the widely used linear reward in the literature (Zhu et al., 2023; Cen et al., 2025).
@@
It holds that $r(x,a)=\phi(x,a)^{\top}\theta^{*}$ where $\phi(x,a):\mathcal{X}\times\mathcal{A}\to\mathbb{R}^{d}$ is the known and fixed feature map, and $\theta^{*}\in\mathbb{R}^{d}$ is the unknown parameter vector. Furthermore, we assume $\lVert\phi(x,a)\rVert_{2}\leq L$ for all $x\in\mathcal{X}$ and $a\in\mathcal{A}$ and $\theta^{*}\in\Theta$ where $\Theta=\{\theta\in\mathbb{R}^{d}\mid\|\theta\|_{2}\leq B\}$.
@@
Then, we can rewrite the probability as $\mathbb{P}\left[y=1\mid x,a,a^{\prime}\right]=\sigma(\phi(x,a)^{\top}\theta^{*}-\phi(x,a^{\prime})^{\top}\theta^{*})$, where $\sigma(w)=\frac{1}{1+\exp(-w)}$. Next, we introduce a key quantity that captures learning complexity.